random231 (random231):

find inverse of the function::: y=x^3+3x+1

random231 (random231):

@RadEn @goformit100 can u guys help me?

OpenStudy (anonymous):

f^-1(x) = 2^(1/3)/[√(x^2 + 2x + 5) - x - 1]^(1/3) - [√(x^2 + 2x + 5) - x - 1]^(1/3)/2^(1/3)

random231 (random231):

how did u get that @cutepochacco ? used wolfram didnt u?

OpenStudy (anonymous):

what you want to do in inverse problems.. if it is y = x something... replace y by x and solve for y. you can do it the other way and sub variables at the end... i always found it easier to just substitute up front and solve for the variable.

random231 (random231):

um @Nameless actually i am used to the process of finding the inverse by taking y. but this is a cubic equation so i am not being able to sepate the x and y. let alone substitution at last....

OpenStudy (anonymous):

http://www.purplemath.com/modules/invrsfcn3.htm Might help :)

random231 (random231):

@PayPay777 when i google search thats the first result shown but its of no help.... i tell u i know the method of finding inverse but it cannot be applied here.

OpenStudy (anonymous):

you are not separating.. just switching around x and y and solving for y. so it will be x=y^3+3y+1. then solve for y.

OpenStudy (anonymous):

basically you will end up with a lot of cubed roots in a rational function. it is kinda nasty in my opinion.

random231 (random231):

yeah its becoming too complicated. the main problem is that this function cannot be made up into any sort of a(y+b)^3+c form.

random231 (random231):

as a result its very hard to sove for y.

random231 (random231):

@Nameless

OpenStudy (anonymous):

i agree.. this one is nasty. a good homework problem... terrible test question. perhaps there is an easier way... but I have not really tried extensively as I am also doing hw

random231 (random231):

yeah this was a test question

random231 (random231):

actually the question was : find the area enclosed by f(x), x=1, x=15, where f(x) is the inverse of the given function. @Nameless !!!!!!!!!

OpenStudy (anonymous):

the given function is the one on the top?

random231 (random231):

yeah

OpenStudy (anonymous):

and bounded by the x axis on the bottom?

random231 (random231):

not given in question

OpenStudy (anonymous):

can you paste the question? are you using calculus to find area? or Algebra?

random231 (random231):

calculus of course

random231 (random231):

area enclosed by y=f(x), x=1,x=15 where f(x) is inverse of g(x)=x^3+3x+1 is a 9/4 b. 18 c 3/4 d. none of these

random231 (random231):

dont say none of these @Nameless lol hahaha

OpenStudy (anonymous):

Here is the inverse generated by Mathematica \[ \frac{\sqrt[3]{\sqrt{x^2-2 x+5}+x-1}}{\sqrt[3]{2}}-\frac{\sqrt[3]{2}}{\sqrt[3]{\sqrt{x^2-2 x+5}+x-1}} \]

random231 (random231):

i know that @eliassaab but i cant figure out how to do it.

OpenStudy (anonymous):

I think it involves finding roots of cubic using formulas similar to the quadratic formula

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cubic_function

ganeshie8 (ganeshie8):

u may use symmetry to avoid finding the inverse

ganeshie8 (ganeshie8):

find the area enclosed by f(x), x=1, x=15, where f(x) is the inverse of the given function. if u think bit hard you can set up the area integral as below :- 2*15 - [ int 0->2 (x^3+3x+1) dx ]

OpenStudy (anonymous):

Agree with @ganeshie8 and the answer is 18

random231 (random231):

yeah ans is 18

random231 (random231):

but why 2*15 i didnt get that!!

random231 (random231):

@ganeshie8 please explain!

ganeshie8 (ganeshie8):

first observation : area between f and y axis = area between inverse(f) and x axis

random231 (random231):

okay

ganeshie8 (ganeshie8):

area between f and y axis in interval y = [1, 15] : int 1 to 15 x(y) dy which is same as : 2*15 - int 0 to 2 y(x)dx

ganeshie8 (ganeshie8):

let me try to show u in graph, it becomes easy to convince :)

random231 (random231):

sure :)

ganeshie8 (ganeshie8):

those two green areas area same, from our first observation of symmetry.

ganeshie8 (ganeshie8):

next we wanto figure out how to compute the area between y = f(x) and y-axis (vertical green region)

ganeshie8 (ganeshie8):

area of vertical green region = (area of rectangle of 2x15) - (area below y=f(x) in that region)

ganeshie8 (ganeshie8):

green + blue = 2*15 blue = 2*15 - blue

random231 (random231):

thanks @ganeshie8 100 medals to u!!! :)